Al is five times as old as Tim. In five years,Al will be four times as old as Tim. How old will Tim be in five years?
Answer by stanbon(75887)

Al is five times as old as Tim. In five years,Al will be four times as old as Tim. How old will Tim be in five years?
-----
Equations:
a = 5t
(a+5) = 4(t+5)
-----
Substitute for "a" and solve for "t":
(5t+5) = 4t+20
t = 15 (Tim's age now)
--------------------------
How old will Tim be in five years?
t+5 = 20 yrs old
=======================
Cheers,
Stan H.
===============